Lee:

Compute the amount of interest earned in the following simple interest problem. A deposit of $1,295 at 7% for 180 days = _____. (Note: Use 365 days in a year) $16,317.00 $444.19 $4.44 $44.42 halp

Lee:

yay

Lee:

how much free time do u have?

justus:

The simple interest formula is \ \[t=(180/365)\] \[P=1,295\] \[r=0.07\]

justus:

simple interest formula is I=P(rt)

Lee:

ok

justus:

now you put the numbers in the formula \

justus:

wait I put numbers but its not working..

justus:

let me retype it

justus:

I=1,295(0.07*0.49)=

justus:

Calculate and get the answer

Lee:

44.4185

justus:

Sorry I took a bit long I had to rewrite my answer 3 times because the equation machine wasnt working

Lee:

k

justus:

Yes but round it up to $44.42

Lee:

thankx

justus:

You're welcome